# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ X Evolution Equations for Density Fluctuations
In light of the preceding analysis, it can be concluded that the quantities represented by $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, are the real physical perturbations to the energy density and particle number density, respectively. Furthermore, it has been demonstrated that the evolution equations for scalar perturbations (37), when combined with the background equations (10), govern the evolution of the energy density perturbation and the particle number density perturbation. In Appendix A, the derivation of evolution equations for their corresponding density fluctuations will be presented. These density fluctuations are defined as follows:
$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{\scriptscriptstyle(0)}(t)},\quad\delta_{n}(t,\bm{x}):=\dfrac{n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{n_{\scriptscriptstyle(0)}(t)}.$ (59)
@@
The final result is a system of evolution equations for the density fluctuations $\delta_{\varepsilon}$ and $\delta_{n}$, in closed, flat, and open flrw universes. These equations are given by
@@
The coefficients $b_{1}$, $b_{2}$ and $b_{3}$ are as follows:
@@
$-\beta^{2}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right),$ (61b)
@@
×$\times\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right).$ (61c)
@@
It is essential to recognize that, with regard to the nonrelativistic limit, the aforementioned equations apply solely in the case of $p\not\equiv 0$.
XI Thermodynamic Quantities
The combined first and second laws of thermodynamics are expressed in terms of density fluctuations, denoted by $\delta_{\varepsilon}$ and $\delta_{n}$ and gauge-invariant expressions are provided for the local pressure and temperature perturbations.
XI.1 Thermodynamic Laws
@@
Eliminating $\mu$ in (63) with the aid of (64) reveals that the combined first and second laws of thermodynamics  (62) can be expressed in a form that does not include $\mu$ or $N$ (see Ref. [33]):
@@
From the background equations (10) and the thermodynamic law (65) it can be demonstrated that $\dot{s}_{\scriptscriptstyle(0)}=0$, indicating that the expansion of the universe takes place without generating entropy. From the transformation (3) it can be concluded that $s_{\scriptscriptstyle(1)}=s_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ is automatically gauge-invariant. By employing (138) and $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$, the thermodynamic relation (65) can be rewritten as
$T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n^{2}_{\scriptscriptstyle(0)}}\left(n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\right).$ (66)
By employing the definitions (59), the following is obtained:
@@
Therefore, $s_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}$ represents the local perturbation to the entropy per particle.
XI.2 Temperature and Pressure Perturbations
Given that both the background temperature, denoted by $T_{{\scriptscriptstyle(0)}}$, and the pressure, represented by $p_{{\scriptscriptstyle(0)}}$, are scalars, their gauge-invariant temperature and pressure perturbations can be defined in a manner analogous to that of the expressions (4):
@@ XI.2 Temperature and Pressure Perturbations
In order to arrive at the gauge-invariant counterpart of (13a), it is first necessary to eliminate $\varepsilon_{{\scriptscriptstyle(1)}}$ and $n_{{\scriptscriptstyle(1)}}$ from (13a) using (4a). The utilization of $\dot{p}_{{\scriptscriptstyle(0)}}=p_{n}\dot{n}_{{\scriptscriptstyle(0)}}+p_{\varepsilon}\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and (68), results in
@@
Eliminating $p_{\varepsilon}$ with the aid of (40) and utilizing the definitions provided in (59) for density fluctuations results in the following expression:
@@
XII Application: Structure Formation in a Flat FLRW Universe
The evolution equations presented in (60), are applicable to open, flat, and closed flrw universes. Henceforth, the discussion will be focused on the flat, ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, universe. Given that the initial density fluctuations occurred in the early universe when $\Lambda\ll\kappa\varepsilon_{\scriptscriptstyle(0)}$, the cosmological constant, $\Lambda$, will be neglected. Consequently, the background equations, as given in Eq. (10), are reduced to
@@ XII Application: Structure Formation in a Flat FLRW Universe
By employing the Friedmann equation (71a), the coefficients (61) of Eq. (60a) result in
@@
XII.1 Era before Decoupling of Matter and Radiation
In this era the primordial fluid is a mixture of radiation and matter, wherein the contribution of matter to the pressure is negligible. Consequently, the equations of state (see, for example, Ref. [8], Eq. (5.49), Ref. [10], Eq. (1.78), and Ref. [17], § V-1) are given by:
@@ XII.1 Era before Decoupling of Matter and Radiation
For the parameter $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and the nonadiabatic speed of sound $\beta$, (40), the following expressions are derived:
@@
The ensuing subsections provide a concise overview of the radiation-dominated and matter-dominated eras preceding decoupling.
XII.1.1 Radiation-dominated Era
@@
In this case one has $\beta^{2}\approx\tfrac{1}{3}$, and $w\approx\tfrac{1}{3}$, implying that $\dot{\beta}\approx 0$ and $p_{\varepsilon}\approx w$. Upon substituting the aforementioned values and expressions (75) into the coefficients (72), the following result is obtained:
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}$
$\qquad\qquad\qquad\;\,\approx\dfrac{-\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\dfrac{\nabla^{2}}{a^{2}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)},$ (78a)
$\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}\approx-\tfrac{3}{4}H\dfrac{n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}.$ (78b)
From (77), it can be deduced that the pressure term on the left-hand side of equation (78a) is, in absolute value, significantly larger than the source term. Consequently, the source term may be neglected. In the radiation-dominated phase, the system of equations (60) can be expressed as follows:
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}\approx 0,$ (79a)
$\left|\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\right|\rightarrow 0.$ (79b)
The limit in Eq. (79b) is a consequence of the fact that the coefficient in Eq. (78b) is negative, since $p_{n}<0$.
In order to solve Eq. (79a), it is necessary to employ the solutions of the background equations (71). These solutions are provided by:
@@
The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{0}\geq 1$. This definition implies that
@@
By employing the Helmholtz equation $\nabla^{2}\delta_{\varepsilon}=-|\bm{q}|^{2}\delta_{\varepsilon}$ and the result derived in (81), it can be shown that the equation (79a) can be rewritten as
$\delta_{\varepsilon}^{\prime\prime}-\dfrac{1}{2\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{\mu_{\text{r}}^{2}}{4\tau}+\dfrac{1}{2\tau^{2}}\right)\delta_{\varepsilon}=0.$ (82)
In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{r}}$ is defined as follows:
@@
In this expression, $\lambda_{0}:=\lambda a(t_{0})$ represents the physical scale of a fluctuation at $t=t_{0}$, with the magnitude of the fluctuation given by $\lvert\bm{q}\rvert=2\pi/\lambda_{0}$.
In order to solve Eq. (82), it is first necessary to replace the variable $\tau$ with the variable defined by $x:=\mu_{\text{r}}\sqrt{\tau}$. Upon transforming back to $\tau$, the general solution of Eq. (82) with constants of integration $A_{1}(\bm{q})$ and $A_{2}(\bm{q})$ is obtained, namely
$\delta_{\varepsilon}(\tau,\bm{q})=\bigl{[}A_{1}(\bm{q})\sin\left(\mu_{\text{r}}\sqrt{\tau}\right)+A_{2}(\bm{q})\cos\left(\mu_{\text{r}}\sqrt{\tau}\right)\bigr{]}\sqrt{\tau},$ (84)
It can be inferred from this solution that, during the radiation-dominated era of the universe, density fluctuations oscillated with an amplitude that grew in a manner proportional to the square root of time.
In the case of large-scale fluctuations, that is, ${\lambda_{0}\rightarrow\infty}$, which leads to $\mu_{\text{r}}\rightarrow 0$, (83), the general solution of Eq. (82) is given by
@@
This solution, with the exception of the precise factors of proportionality, has been derived by a significant number of authors, as evidenced in Refs. [2, 39, 40, 41, 42].
A comparison of the solutions (85) and (155a) reveals that the solution ${\propto\tau}$ is a consequence of the homogeneous part of equation (154a). The solution $\propto\tau^{1/2}$ is the particular solution of the inhomogeneous equation. This solution can only be obtained if the covariant divergence, $\vartheta_{{\scriptscriptstyle(1)}}$, is taken into account, as demonstrated in Appendix C.1. The minus sign preceding the solution $\propto\tau$ is a direct consequence of the fact that ${\vartheta_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}>0}$, as follows from (155b). The presence of the gauge function $\psi$ precludes the derivation of the solution (85) from (154). As demonstrated in Appendix A, $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the left-hand side of (78a), as evidenced by the derivation of (60a).
XII.1.2 Matter-dominated Era
The radiation-dominated phase came to an end and the matter-dominated phase commenced at $z(t_{\text{eq}})=3387$, as indicated in Table 1, well before decoupling [43], when
$a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}=n_{{\scriptscriptstyle(0)}}mc^{2}.$ (86)
Given that $p_{n}<0$, it can be shown from Eq. (60b) that
It can thus be seen that fluctuations in the energy density are coupled to those in the particle number density. In light of the fact that the evolution equations (60) are general relativistic, it follows that the weak equivalence principle is valid. This implies that the evolution of fluctuations in the particle number density is independent of the composition of matter. Therefore, if cdm consists of particles that interact via gravitation, the fluctuations in the cdm are also linked to the fluctuations in the energy density. This implies that it is impossible for cdm to have contracted before decoupling. Consequently, cdm cannot have initiated the formation of structure after decoupling.
XII.2 Era after Decoupling of Matter and Radiation
Once protons and electrons have combined to form hydrogen, the radiation pressure will be negligible, and the equations of state will be those of a nonrelativistic monatomic perfect gas with three degrees of freedom. These are given by the equations of state [see Ref. [31], Eqs. (15.8.20)–(15.8.21), and Ref. [33], Eq. (13)]:

### v3

@@ XI Relativistic Cosmological Perturbation Theory
As demonstrated in Sec. VIII, the evolution equations for scalar perturbations (37) when combined with the background equations (11) describe the evolution of energy density and particle number density perturbations. In Appendix A, the derivation of evolution equations for their corresponding relative density perturbations defined by
$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{\scriptscriptstyle(0)}(t)},\quad\delta_{n}(t,\bm{x}):=\dfrac{n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{n_{\scriptscriptstyle(0)}(t)},$ (61)
@@
will be presented. The result is a system of evolution equations for the relative density perturbations $\delta_{\varepsilon}$ and $\delta_{n}$, in closed, flat, and open flrw universes, given by
@@
The background equations (11) and an equation of state (2) for the pressure determine the coefficients $b_{1}$, $b_{2}$, and $b_{3}$. These coefficients are given by
@@
[…]
@@
[…]
@@
The local curvature perturbation, ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, caused by a local density perturbation as well as the divergence, ${\vartheta}_{{\scriptscriptstyle(1)}}$, of the local fluid velocity caused by local pressure gradients, are incorporated into system (62) via expressions (127) and (132), respectively.
The system of equations (62) allows us to study the evolution of the relative density perturbations (61) in closed $K=+1$, flat $K=0$, and open K=1K=-1 flrw universes, where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=-6K/a^{2}$.
Sec. XIII will demonstrate that $p_{n}<0$ before and after the decoupling of radiation and matter. For $p_{n}>0$, however, the coefficient in Eq. (62b) is positive, causing $\delta_{n}-\delta_{\varepsilon}/(1+w)$ to grow exponentially, which is not physically plausible.
In the system (62), Eq. (62a) is the evolution equation for energy density perturbations. Since $p_{n}<0$ for an equation of state $p=p(n,\varepsilon)$, the right-hand side of Eq. (62a) is nonzero. This indicates that density perturbations are nonadiabatic. Density perturbations are adiabatic when the source term of Eq. (62a) vanishes, i.e., when
$\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}=0.$ (64)
Therefore, Eq. (62b) can be regarded as the entropy equation. An explanation is provided by Sec. XII.1.
Since $p_{n}<0$, the entropy equation (62b) shows that
This means that relative perturbations in energy density are coupled with those in particle number density. Since general relativity is indifferent to the composition of matter, this is true for cdm as well as for ordinary matter, where it is assumed that cdm interacts only via gravity. Therefore, the contraction of cdm before decoupling was impossible. This precludes the formation of potential wells into which ordinary matter could fall after decoupling to form structures. Consequently, the initiation of structure formation by cdm after decoupling is also impossible.
XII Thermodynamic Quantities
To link an equation of state to Eqs. (62), one needs gauge-invariant expressions for the entropy, temperature, and pressure perturbations. These quantities must be expressed in terms of the relative density perturbations $\delta_{\varepsilon}$ and $\delta_{n}$. This section presents the derivation of these expressions.
XII.1 Entropy Perturbations
@@
Eliminating $\mu$ in (67) with the aid of (68) reveals that the combined first and second laws of thermodynamics (66) can be expressed in a form that does not include $\mu$ and $N$ (see Ref. [32]):
@@
From the background equations (11) and the thermodynamic law (69) it can be demonstrated that $\dot{s}_{\scriptscriptstyle(0)}=0$}, indicating that the expansion of the universe takes place without generating entropy. From the definition of gauge-invariant quantities (7), it can be concluded that $s_{\scriptscriptstyle(1)}=s_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ is automatically gauge-invariant. By using (136) and $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$, the thermodynamic relation (69) can be rewritten as
[…]
By using the definitions (61), the following is obtained:
@@
Therefore, Eq. (62b) can be regarded as the evolution equation for the entropy per particle.
XII.2 Temperature and Pressure Perturbations
Since the background temperature $T_{{\scriptscriptstyle(0)}}$ and pressure $p_{{\scriptscriptstyle(0)}}$ are scalars, their gauge-invariant perturbations are defined using the definition (7):
@@ XII.2 Temperature and Pressure Perturbations
The expression for $T^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ will be used in Sec. XIII.2. In order to arrive at the gauge-invariant counterpart of (14a), it is first necessary to eliminate $\varepsilon_{{\scriptscriptstyle(1)}}$ and $n_{{\scriptscriptstyle(1)}}$ from (14a) using (6a). Using $\dot{p}_{{\scriptscriptstyle(0)}}=p_{n}\dot{n}_{{\scriptscriptstyle(0)}}+p_{\varepsilon}\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and (72), yields
@@
Eliminating $p_{\varepsilon}$ with the aid of (40) and using the definitions provided in (61) for relative density perturbations results in the following expression:
@@
XIII Application: Flat FLRW Universe
The evolution equations presented in (62), are applicable to open, flat, and closed flrw universes. We will now focus on studying density perturbations in a flat flrw universe, for which ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$.
Given that the initial density perturbations occurred in the early universe when $\Lambda\ll\kappa\varepsilon_{\scriptscriptstyle(0)}$, the cosmological constant $\Lambda$ will be neglected. Consequently, the background equations (11), are reduced to
@@ XIII Application: Flat FLRW Universe
By using the Friedmann equation (75a), the coefficients (63) of Eq. (62a) result in
@@
XIII.1 Era before Decoupling of Matter and Radiation
In this era the primordial fluid is a mixture of radiation and matter, wherein the contribution of matter to the pressure is negligible. Consequently, the equations of state are given by [see Ref. [13], § V-1]:
@@ XIII.1 Era before Decoupling of Matter and Radiation
Since $p_{n}<0$, energy density perturbations are coupled to particle number density perturbations according to (65).
For the parameter $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and the nonadiabatic speed of sound $\beta$, (40), it is found
@@
In this case one has $\beta^{2}\approx\tfrac{1}{3}$, and $w\approx\tfrac{1}{3}$, implying that $\dot{\beta}\approx 0$. Upon substituting these values and expressions (79) into the coefficients (76), and using (81) the result is
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}=0.$ (82)
In order to solve Eq. (82), it is necessary to use the solutions to the background equations (75). These solutions are given by
@@
The dimensionless time, denoted by the symbol $\tau$, is defined by $\tau:=t/t_{0}\geq 1$. This definition implies that
@@
By using the Helmholtz equation $\nabla^{2}\delta_{\varepsilon}=-|\bm{q}|^{2}\delta_{\varepsilon}$ and the result derived in (84), it can be shown that the equation (82) can be rewritten as
$\delta_{\varepsilon}^{\prime\prime}-\dfrac{1}{2\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{\mu_{\text{r}}^{2}}{4\tau}+\dfrac{1}{2\tau^{2}}\right)\delta_{\varepsilon}=0,$ (85)
where the prime denotes differentiation with respect to the $\tau$. With $|\bm{q}|=2\pi/\lambda_{0}$, the parameter $\mu_{\text{r}}$ is defined by
@@
In this expression, $\lambda_{0}:=\lambda a(t_{0})$ represents the physical scale of a density perturbation at $t=t_{0}$.
In order to solve Eq. (85), it is first necessary to replace the variable $\tau$ with the variable defined by $x:=\mu_{\text{r}}\sqrt{\tau}$. Upon transforming back to $\tau$, the general solution of Eq. (85) with constants of integration $A_{1}(\bm{q})$ and $A_{2}(\bm{q})$ is obtained:
$\delta_{\varepsilon}(\tau,\bm{q})=\bigl[A_{1}(\bm{q})\sin\left(\mu_{\text{r}}\sqrt{\tau}\right)+A_{2}(\bm{q})\cos\left(\mu_{\text{r}}\sqrt{\tau}\right)\bigr]\sqrt{\tau}.$ (87)
Since $\vartheta_{{\scriptscriptstyle(1)}}$ is included in Eq. (82) it yields oscillating relative density perturbations with an increasing amplitude. In contrast, the standard equation (153), for which $\vartheta_{{\scriptscriptstyle(1)}}=0$ (i.e., no fluid flow) yields oscillating solutions with a constant amplitude.
In the case of large-scale perturbations, that is, $\lambda_{0}\rightarrow\infty$}, which leads to $\mu_{\text{r}}\rightarrow 0$, (86), the general solution to Eq. (85) is given by
@@
A comparison of the solutions (88) and (155a) reveals that the solution proportional to $\tau$ in (88) is also a solution to the homogeneous part of equation (154a). The solution proportional to $\tau^{1/2}$ is the particular solution to the inhomogeneous equation. This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.1. The minus sign preceding the solution proportional to $\tau$ is a consequence of the fact that $\vartheta_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}>0$}, as follows from (155b). The presence of the gauge function $\psi(\bm{x})$ precludes the derivation of the solution (88) from (154).
XIII.2 Era after Decoupling of Matter and Radiation
Once protons and electrons have combined to form hydrogen, radiation pressure will be negligible and the equations of state will be those of a nonrelativistic, monatomic, perfect gas with three degrees of freedom. Therefore, pressure can be neglected in the background universe. However, the momentum conservation law (49) indicates that, in order to study the evolution of density perturbations, one must consider pressure perturbations, since density perturbations are static (51) in a pressureless fluid. According to the literature [see Ref. [32], Eq. (13), and Ref. [42], Eqs. (15.8.20)–(15.8.21)], we use the following equations of state for a nonrelativistic gas: